The average grade/mark of all college students is 70 with standard deviation of 20. If a random sample of 49 students is taken, what is the probability that the sample average grade/mark is
a) At least 65
b) Less than 78
c) Between 65 and 72
1
Expert's answer
2020-05-18T03:56:56-0400

"P(65\le)"

"Z=\frac{65-70}{\frac{20}{7}}=-1.75"

"P(65\le)= 1-P(Z\le-1.75)"

From z table, we have "1-0.040059=0.959941"

"P(78\ge)"

"Z=\frac{78-70}{\frac{20}{7}}=2.8"

"P(78\ge)= P(Z\le2.8)"

From z table, we have 0.997445

"P(65\le x\le72)"

"Z=\frac{72-70}{\frac{20}{7}}=0.7"

"P(65\le x\le72)=P(-1.75\le Z\le0.7)"

From z tables we have "0.758036-0.040059=0.717977"